# RB-SCD: A New Benchmark for Semantic Change Detection of Roads and Bridges in Traffic Scenes

### v1

@@ 3.1 Data Annotation and Analysis
Diverse Geographic Coverage: As shown in Figure 4, following the principles in (Long et al., 2021), we collect 260 pairs of RS images with resolutions ranging from 270 × 534 to 7,215 × 4,366 (0.59 m/pixel), which can be cropped into 5,016 pairs of 256 × 256 patches, including 4,015 pairs for the training set and 1,001 pairs for the testing set. The dataset integrates satellite images obtained primarily from Google Earth, covering various regions worldwide, including Chinese cities such as Hefei, Xi’an, Chongqing and Guangzhou, as well as representative areas from Japan, Singapore, Nigeria, Italy, America, and Indonesia, etc. This extensive geographic coverage ensures that the dataset encompasses traffic scenes under varying climatic conditions, geological structures, and urban planning contexts, thereby guaranteeing its diversity and representativeness.
@@
Multiple Road and Bridge Types: As shown in Figure 5, the dataset encompasses various types of roads and bridges from different geographical regions and urban environments. The road categories include concrete and asphalt roads used for urban main roads, secondary roads, residential streets, and highways, comprehensively reflecting the common transportation routes found in cities. Additionally, the dataset contains elevated roads designed to alleviate congestion and improve traffic efficiency, which are typically found in densely populated or high-traffic urban areas. In terms of bridge structures, the dataset includes large cable-stayed bridges that span wide rivers or deep valleys, as well as girder bridges and pedestrian bridges used for short- to medium-span connections and pedestrian access. The dataset also includes other types of specialized roads and bridges, demonstrating a variety of engineering styles and applications. This richness in road and bridge types enables the dataset to represent complex traffic environments and provides robust support for training CD models across multiple scales and diverse scenarios.
3.2 Label Imbalance Analysis and Challenges
@@
As shown in Figure  3, our RB-SCD dataset exhibits a certain degree of sample imbalance, primarily manifested in the unequal distribution of different road and bridge types as well as change categories. Due to the diversity in data collection times and geographical regions, some categories of roads, bridges, and change types have fewer samples. In terms of semantic changes, there is also an imbalance in the types of changes in the dataset, especially for fine-grained CD tasks. For some types of changes (e.g., road widening, bridge construction), there are more samples, while for other types of changes (e.g., road repair, demolition), there are fewer samples. This imbalance may lead to the model overfitting to the more common change types, while neglecting the rarer ones. Additionally, the imbalance between change and no-change samples is another issue, particularly for CD tasks. The no-change samples are relatively more abundant, which may lead to a decrease in the model’s sensitivity when processing change samples. To improve the accuracy and robustness of the model, we will further collect more samples of different types of changes in future work and employ techniques such as data augmentation and sample weighting to alleviate the model’s difficulty in detecting rare changes.
4 Methodology
4.1 Overview
@@
As shown in Figure 6, our MFDCD mainly includes two components: (1) DFC module effectively couples high-level and low-level visual features with high-frequency and low-frequency spectral features. (2) TFF module enhances the text information obtained through zero-shot inference, explores the representation of text features in the frequency domain, and promotes the interaction and fusion of multimodal data. The details are as follows.
4.2 DFC Module
@@
Given the bi-temporal images $T_{1}$ and $T_{2}$, ResNet50 (He et al., 2016) is used as the backbone network to extract the corresponding multi-scale features $F_{i}^{t}(i\in\left\{1,2,3,4\right\},t\in\left\{1,2)\right\}$. The features from the first two layers of the network are considered as low-level features, while the features from the last two layers are considered as high-level features. The low-frequency and high-frequency components of these features are then obtained through wavelet transform, which allows for the separation of coarse and fine-grained information in both spatial and frequency domains.
@@
Adaptive Sparse Frequency Fusion: Low-level features are rich in detailed information, while high-frequency features contain information about edges, details, and abrupt changes, which naturally correspond to low-level features. However, high-frequency features are sensitive to noise and may amplify the influence of irrelevant regions. As shown in the upper-right corner of Figure  6, the ASFF block applies a novel sparsity constraint to the high-frequency features, which is dynamically adjusted during training iterations. By combining the sparsity-constrained high-frequency features with low-level features, both the image’s detailed information (such as edges and textures) and local structural information are preserved. This combination makes the network more responsive to CD, especially in areas with detailed changes. The process is as follows:
@@
Bidirectional Temporal Frequency Fusion: As shown in the lower-right corner of Figure 6. Take the low-frequency features of time $T_{2}$ and the high-level features of time $T_{1}$ as an example, high-level features at time $T_{1}$ typically contain the semantic hierarchy of the image, enabling the expression of global structural changes in the image, such as road widening and bridge construction. Low-frequency features at time $T_{2}$ mainly focus on smooth regions or large-scale change information in the image. By combining high-level features from $T_{1}$ with low-frequency features from $T_{2}$, the network can capture global structural changes across time points while maintaining sensitivity to large-scale, smooth region changes. This models the bidirectional relationship of changes: local changes at $T_{1}$ may lead to semantic-level changes at $T_{2}$, while detailed changes at $T_{2}$ may arise from regional distribution adjustments at $T_{1}$. The details of BTFF block are as follows:
@@
4.3 TFF Module
As shown in Figure 6, we leverage the powerful zero-shot inference capability of the CLIP model to obtain textual information from RS images for multimodal modeling. For the categories inferred by the CLIP model, we select the top five categories with the highest probability values. For the foreground features, the text description $T_{d}$ is designed in the format “Remote sensing image at time $\left\{period\right\}$ has a $\left\{value\right\}$ probability of being the $\left\{key\right\}$ ," where $\left\{period\right\}$ represents either $T_{1}$ or $T_{2}$, $\left\{value\right\}$ is the probability value of the corresponding category and $\left\{key\right\}$ represents the category itself. $E_{T}$ is a pre-trained CLIP text encoder that encodes the text embedding. Considering that discrete Fourier transform can help to reveal high-order semantic patterns in text, especially by identifying more complex semantic information through the descriptors used at different temporal points ($T_{1}$, $T_{2}$). Through Fourier transform, periodic or repetitive structures in the text can be extracted, which helps identify commonalities and differences in changes. We first use Fourier transform to convert the text embeddings from the text domain to the frequency domain:
@@ 4.3 TFF Module
where $\mathfrak{R}(\mathbf{T}_{freq})$ and $\mathfrak{I}(\mathbf{T}_{freq})$ represent the real and imaginary parts of the frequency domain features, respectively. $\mathbf{E}^{(k)}$ represents the node features after the $k$-th filter, $\hat{\mathbf{A}}$ is the normalized adjacency matrix, which represents the relationships between nodes, $\mathbf{W}_{k}$ is the weight matrix of the graph convolution filter at the $k$-th layer, and $\sigma$ is the activation function, $\mathbf{E}^{f}$ is the weighted sum of the outputs from multiple filters. We use the text-image encoder $E_{VT}$ of the CLIP model to fuse $\mathbf{E}^{f}$ and the image features. After subtracting the results from the bi-temporal phases, the output is fed into the decoder (Chen et al., 2024) to obtain our CD results.
4.4 Frequency Transform Function Analysis
Why use discrete wavelet transform for image features and discrete Fourier transform for textual features? The advantage of wavelet transform lies in its ability to analyze images at different scales and locations, making it particularly suitable for multi-level, multi-scale image analysis tasks. As shown in Figure  2, in remote sensing image CD, wavelet transform helps to extract high-frequency features (e.g., detailed changes) and low-frequency features (e.g., overall structure), aiding in the capture of changes at different spatiotemporal scales in the image. Considering that text is typically discrete sequential data, its changes are often global or frequency-related. Therefore, on the other hand, the Fourier transform can effectively reveal periodic semantic patterns in the text, which contributes to identify patterns and trends in text changes.
5 Experiments and Results
5.1 Public CD Datasets
In addition to using our proposed dataset, as shown in Table 3, we have employed three public CD datasets to evaluate the robustness of our model. The LEVIR-CD (Chen and Shi, 2020) dataset consists of 637 pairs of RS images with a resolution of 1,024 × 1,024. It mainly focuses on building area changes, such as building growth. The WHU-CD (Ji et al., 2018) dataset comprises a pair of high-resolution RS images with a size of 32,507 × 15,354 pixels. It covers areas before and after an earthquake and primarily includes building renovations and additions. The SYSU-CD (Shi et al., 2021) dataset is a collection of 20,000 pairs of high-resolution aerial images, each with a size of 256 × 256 pixels. The images are captured in the Hong Kong area and include urban construction, suburban expansion, and vegetation changes, etc.
Table 3: Public CD datasets used in the experiments.
@@ 5.1 Public CD Datasets
SYSU-CD 12000/ 40000/ 4000 Aerial images 0.5 m/pixel
@@
Table 4: Comparison of quantitative performance for semantic change detection on RB-SCD dataset. All metrics are expressed as percentages (%). Results highlighted in red indicate the best performance, while those in blue denote the second-best performance.
Class HGINet (Long et al., 2024) CdSC (Wang et al., 2024) DEFO (Li et al., 2024) ChangeCLIP (Dong et al., 2024) Ours
@@
Table 5: Performance of the RS foundation model LWGANet on RB-SCD dataset. All metrics are expressed as percentages (%).
Class LWGANet-L1 (Lu et al., 2025) LWGANet-L2 (Lu et al., 2025)
@@
background 98.3 99.14 98.84 99.44 98.27 99.13 98.8 99.45
bare land ->road 54.17 70.27 73.28 67.5 53.17 69.42 74.46 65.02
@@
road->vegetation 15.54 26.89 55.44 17.75 13.7 24.1 68.9 14.6
vegetation->road 59.14 74.32 75.55 73.13 60.1 75.08 78.81 71.68
building->road 42.1 59.25 81.97 46.4 42.27 59.42 76.5 48.57
@@
Table 6: Comparison of quantitative performance for binary change detection on RB-SCD dataset. All metrics are expressed as percentages (%). Results highlighted in red indicate the best performance, while those in blue denote the second-best performance.
@@
FC-diff (Caye Daudt et al., 2018) 57.81 73.26 69.63 77.30
FC-ef (Caye Daudt et al., 2018) 59.05 74.26 69.05 80.31
FC-cat (Caye Daudt et al., 2018) 60.34 75.26 71.99 78.85
SNUNet (Fang et al., 2022) 62.90 77.22 72.47 82.64
BIT (Chen et al., 2022) 62.21 76.71 70.97 83.69
ChangeFormer (Bandara and Patel, 2022) 61.88 76.45 69.69 84.66
ICIF (Feng et al., 2022) 61.80 76.39 71.93 81.43
TFI-GR (Li et al., 2022) 64.72 78.58 73.78 84.05
DMINet (Feng et al., 2023) 62.99 77.29 72.14 83.24
USSFC-Net (Lei et al., 2023) 62.55 76.96 75.07 78.95
ELGC-Net (Noman et al., 2024) 65.53 79.18 73.99 85.15
@@
Table 7: Comparison of quantitative performance of various CD techniques on public datasets. All metrics are expressed as percentages (%). Results highlighted in red indicate the best performance, while those in blue denote the second-best performance. Results marked with * are obtained from the original papers.
Model LEVIR-CD (Chen and Shi, 2020) SYSU-CD (Shi et al., 2021) WHU-CD (Ji et al., 2018)
@@
FC-diff (Caye Daudt et al., 2018) 75.18 85.83 84.59 87.12 42.05 59.20 46.72 80.80 69.48 81.99 78.90 89.14
FC-ef (Caye Daudt et al., 2018) 71.80 83.59 82.27 84.94 55.83 71.65 65.41 79.22 58.53 71.65 65.41 79.22
FC-cat (Caye Daudt et al., 2018) 76.98 86.99 85.64 88.38 60.68 75.53 71.38 80.21 72.24 83.87 84.22 83.53
SNUNet (Fang et al., 2022) 82.20 90.23 88.97 91.54 66.73 80.04 79.79 80.30 76.02 86.38 88.95 83.95
BIT (Chen et al., 2022) 82.06 90.15 88.98 91.34 62.50 76.93 72.50 81.93 80.40 89.13 89.69 88.58
ChangeFormer (Bandara and Patel, 2022) 81.93 90.07 87.38 92.92 60.46 75.36 71.06 80.22 83.88 91.23 90.76 91.71
ICIF (Feng et al., 2022) 79.86 88.80 86.85 90.84 61.38 76.07 73.31 91.94 78.27 87.81 84.06 91.96
TFI-GR (Li et al., 2022) 83.03 90.73 89.53 91.95 71.38 83.30 81.16 85.56 88.55 93.92 92.65 95.24
DMINet (Feng et al., 2023) 83.14 90.79 88.89 92.78 68.20 81.10 76.82 85.88 79.53 88.60 91.24 86.10
USSFC-Net (Lei et al., 2023) 82.54 90.44 91.82 89.09 65.75 79.34 81.52 77.27 83.92 91.26 93.20 89.40
M-Swin * (Pan et al., 2024) 83.57 91.05 90.05 92.07 -   -   -   - 86.09 92.94 91.37 94.56
ELGC-Net (Noman et al., 2024) 83.52 91.02 89.78 92.30 65.34 79.03 77.26 80.89 75.56 86.08 82.15 90.40
MDENet * (Liu et al., 2024) 83.63 91.0992.68 89.55 -   -   -   - 87.74 93.47 95.49 91.54
@@
5.2 Implementation Details
Our experiments are conducted on a single NVIDIA GeForce RTX 3090 GPU. The AdamW (Kingma and Ba, 2014) is employed, with a learning rate of 0.0001 and a weight decay parameter of 0.01. Cross entropy loss is used as our loss function, and the maximum number of iterations is set to 40,000. For data augmentation, RandomFlip and PhotoMetricDistortion are applied. The constant $g$ in the ASFF block is set to 0.1, the decay rate $\gamma$ is set to 0.0001, and $\lambda_{0}$ is set to 1. To evaluate the performance of our experiments, we adopt four widely used metrics: recall (Rec), precision (Pre), F1-score (F1) and intersection over union (IoU), where IoU and F1 are the primary metrics, while the others are auxiliary metrics. The following are formulas for defining the evaluation metrics as mentioned above.

### v2

@@ III-A Data Annotation and Analysis
Diverse Geographic Coverage: As shown in Fig. 4, following the principles in [44], we collect 260 pairs of RS images with resolutions ranging from 270 × 534 to 7,215 × 4,366 (0.59 m/pixel), which can be cropped into 5,016 pairs of 256 × 256 patches, including 4,015 pairs for the training set and 1,001 pairs for the testing set. The dataset integrates satellite images obtained primarily from Google Earth, covering various regions worldwide, including Chinese cities such as Hefei, Xi’an, Chongqing and Guangzhou, as well as representative areas from Japan, Singapore, Nigeria, Italy, America, and Indonesia, etc. This extensive geographic coverage ensures that the dataset encompasses traffic scenes under varying climatic conditions, geological structures, and urban planning contexts, thereby guaranteeing its diversity and representativeness.
@@
Multiple Road and Bridge Types: As shown in Fig. 5, the dataset encompasses various types of roads and bridges from different geographical regions and urban environments. The road categories include concrete and asphalt roads used for urban main roads, secondary roads, residential streets, and highways, comprehensively reflecting the common transportation routes found in cities. Additionally, the dataset contains elevated roads designed to alleviate congestion and improve traffic efficiency, which are typically found in densely populated or high-traffic urban areas. In terms of bridge structures, the dataset includes large cable-stayed bridges that span wide rivers or deep valleys, as well as girder bridges and pedestrian bridges used for short- to medium-span connections and pedestrian access. The dataset also includes other types of specialized roads and bridges, demonstrating a variety of engineering styles and applications. This richness in road and bridge types enables the dataset to represent complex traffic environments and provides robust support for training CD models across multiple scales and diverse scenarios.
III-B Label Imbalance Analysis and Challenges
@@
As shown in Fig. 3, our RB-SCD dataset exhibits a certain degree of sample imbalance In terms of semantic changes, there is an imbalance in the types of changes in the dataset, especially for fine-grained CD tasks. For some types of changes (e.g., road widening, bridge construction), there are more samples, while for other types of changes (e.g., road repair, demolition), there are fewer samples. This imbalance may lead to the model overfitting to the more common change types, while neglecting the rarer ones. Additionally, the imbalance between change and no-change samples is another issue, particularly for CD tasks. The no-change samples are relatively more abundant, which may lead to a decrease in the model’s sensitivity when processing change samples. To improve the accuracy and robustness of the model, we will further collect more samples of different types of changes in future work and employ techniques such as data augmentation and sample weighting to alleviate the model’s difficulty in detecting rare changes.
IV Methodology
IV-A Overview
@@
As shown in Fig. 6, our MFDCD mainly includes two components: (1) DFC module effectively couples high-level and low-level visual features with high-frequency and low-frequency spectral features. (2) TFF module enhances the text information obtained through zero-shot inference, explores the representation of text features in the frequency domain, and promotes the interaction and fusion of multimodal data. The details are as follows.
IV-B DFC Module
@@
Given the bi-temporal images $T_{1}$ and $T_{2}$, ResNet50 [45] is used as the backbone network to extract the corresponding multi-scale features $F_{i}^{t}(i\in\left\{1,2,3,4\right\},t\in\left\{1,2)\right\}$. The features from the first two layers of the network are considered as low-level features, while the features from the last two layers are considered as high-level features. The low-frequency and high-frequency components of these features are then obtained through wavelet transform, which allows for the separation of coarse and fine-grained information in both spatial and frequency domains.
@@
Adaptive Sparse Frequency Fusion: Low-level features are rich in detailed information, while high-frequency features contain information about edges, details, and abrupt changes, which naturally correspond to low-level features. However, high-frequency features are sensitive to noise and may amplify the influence of irrelevant regions. As shown in the upper-right corner of Fig. 6, the ASFF block applies a novel sparsity constraint to the high-frequency features, which is dynamically adjusted during training iterations. By combining the sparsity-constrained high-frequency features with low-level features, both the image’s detailed information (such as edges and textures) and local structural information are preserved. This combination makes the network more responsive to CD, especially in areas with detailed changes. The process is as follows:
@@
Bidirectional Temporal Frequency Fusion: As shown in the lower-right corner of Fig. 6. Take the low-frequency features of time $T_{2}$ and the high-level features of time $T_{1}$ as an example, high-level features at time $T_{1}$ typically contain the semantic hierarchy of the image, enabling the expression of global structural changes in the image, such as road widening and bridge construction. Low-frequency features at time $T_{2}$ mainly focus on smooth regions or large-scale change information in the image. By combining high-level features from $T_{1}$ with low-frequency features from $T_{2}$, the network can capture global structural changes across time points while maintaining sensitivity to large-scale, smooth region changes. This models the bidirectional relationship of changes: local changes at $T_{1}$ may lead to semantic-level changes at $T_{2}$, while detailed changes at $T_{2}$ may arise from regional distribution adjustments at $T_{1}$. The details of BTFF block are as follows:
@@
IV-C TFF Module
As shown in Fig. 6, we leverage the powerful zero-shot inference capability of the CLIP model to obtain textual information from RS images for multimodal modeling. For the categories inferred by the CLIP model, we select the top five categories with the highest probability values. For the foreground features, the text description $T_{d}$ is designed in the format “Remote sensing image at time $\left\{period\right\}$ has a $\left\{value\right\}$ probability of being the $\left\{key\right\}$ ,” where $\left\{period\right\}$ represents either $T_{1}$ or $T_{2}$, $\left\{value\right\}$ is the probability value of the corresponding category and $\left\{key\right\}$ represents the category itself. $E_{T}$ is a pre-trained CLIP text encoder that encodes the text embedding. Considering that discrete Fourier transform can help to reveal high-order semantic patterns in text, especially by identifying more complex semantic information through the descriptors used at different temporal points ($T_{1}$, $T_{2}$). Through Fourier transform, periodic or repetitive structures in the text can be extracted, which helps identify commonalities and differences in changes. We first use Fourier transform to convert the text embeddings from the text domain to the frequency domain:
@@ IV-C TFF Module
where $\mathfrak{R}(\mathbf{T}_{freq})$ and $\mathfrak{I}(\mathbf{T}_{freq})$ represent the real and imaginary parts of the frequency domain features, respectively. $\mathbf{E}^{(k)}$ represents the node features after the $k$-th filter, $\hat{\mathbf{A}}$ is the normalized adjacency matrix, which represents the relationships between nodes, $\mathbf{W}_{k}$ is the weight matrix of the graph convolution filter at the $k$-th layer, and $\sigma$ is the activation function, $\mathbf{E}^{f}$ is the weighted sum of the outputs from multiple filters. We use the text-image encoder $E_{VT}$ of the CLIP model to fuse $\mathbf{E}^{f}$ and the image features. After subtracting the results from the bi-temporal phases, the output is fed into the decoder [46] to obtain our CD results.
IV-D Frequency Transform Function Analysis
Why use discrete wavelet transform for image features and discrete Fourier transform for textual features? The advantage of wavelet transform lies in its ability to analyze images at different scales and locations, making it particularly suitable for multi-level, multi-scale image analysis tasks. As shown in Fig. 2, in remote sensing image CD, wavelet transform helps to extract high-frequency features (e.g., detailed changes) and low-frequency features (e.g., overall structure), aiding in the capture of changes at different spatiotemporal scales in the image. Considering that text is typically discrete sequential data, its changes are often global or frequency-related. Therefore, on the other hand, the Fourier transform can effectively reveal periodic semantic patterns in the text, which contributes to identify patterns and trends in text changes.
V Experiments and Results
V-A Public CD Datasets
In addition to using our proposed dataset, as shown in Tab. III, we have employed three public CD datasets to evaluate the robustness of our model. The LEVIR-CD [7] dataset consists of 637 pairs of RS images with a resolution of 1,024 × 1,024. It mainly focuses on building area changes, such as building growth. The WHU-CD [6] dataset comprises a pair of high-resolution RS images with a size of 32,507 × 15,354 pixels. It covers areas before and after an earthquake and primarily includes building renovations and additions. The SYSU-CD [8] dataset is a collection of 20,000 pairs of high-resolution aerial images, each with a size of 256 × 256 pixels. The images are captured in the Hong Kong area and include urban construction, suburban expansion, and vegetation changes, etc.
TABLE III: Public CD datasets used in the experiments.
@@ V-A Public CD Datasets
SYSU-CD 12000/ 4000/ 4000 Aerial images 0.5 m/pixel
@@
TABLE IV: Comparison of quantitative performance for semantic change detection on RB-SCD dataset. All metrics are expressed as percentages (%). Results highlighted in red indicate the best performance, while those in blue denote the second-best performance.
Class HGINet [37] CdSC [36] DEFO [35] ChangeCLIP [23] Ours
@@
TABLE V: Performance of the RS foundation model LWGANet on RB-SCD dataset. All metrics are expressed as percentages (%).
Class LWGANet-L1 [47] LWGANet-L2 [47]
@@
background 98.30 99.14 98.84 99.44 98.27 99.13 98.8 99.45
bare land ->road 54.17 70.27 73.28 67.50 53.17 69.42 74.46 65.02
@@
road->vegetation 15.54 26.89 55.44 17.75 13.70 24.10 68.9 14.60
vegetation->road 59.14 74.32 75.55 73.13 60.10 75.08 78.81 71.68
building->road 42.10 59.25 81.97 46.40 42.27 59.42 76.50 48.57
@@
TABLE VI: Comparison of quantitative performance for binary change detection on RB-SCD dataset. All metrics are expressed as percentages (%). Results highlighted in red indicate the best performance, while those in blue denote the second-best performance.
@@
FC-diff [15] 57.81 73.26 69.63 77.30
FC-ef [15] 59.05 74.26 69.05 80.31
FC-cat [15] 60.34 75.26 71.99 78.85
SNUNet [16] 62.90 77.22 72.47 82.64
BIT [18] 62.21 76.71 70.97 83.69
ChangeFormer [19] 61.88 76.45 69.69 84.66
ICIF [17] 61.80 76.39 71.93 81.43
TFI-GR [48] 64.72 78.58 73.78 84.05
DMINet [20] 62.99 77.29 72.14 83.24
USSFC-Net [49] 62.55 76.96 75.07 78.95
ELGC-Net [50] 65.53 79.18 73.99 85.15
@@
TABLE VII: Comparison of quantitative performance of various CD techniques on public datasets. All metrics are expressed as percentages (%). Results highlighted in red indicate the best performance, while those in blue denote the second-best performance. Results marked with * are obtained from the original papers.
Model LEVIR-CD [7] SYSU-CD [8] WHU-CD [6]
@@
FC-diff [15] 75.18 85.83 84.59 87.12 42.05 59.20 46.72 80.80 69.48 81.99 78.90 89.14
FC-ef [15] 71.80 83.59 82.27 84.94 55.83 71.65 65.41 79.22 58.53 71.65 65.41 79.22
FC-cat [15] 76.98 86.99 85.64 88.38 60.68 75.53 71.38 80.21 72.24 83.87 84.22 83.53
SNUNet [16] 82.20 90.23 88.97 91.54 66.73 80.04 79.79 80.30 76.02 86.38 88.95 83.95
BIT [18] 82.06 90.15 88.98 91.34 62.50 76.93 72.50 81.93 80.40 89.13 89.69 88.58
ChangeFormer [19] 81.93 90.07 87.38 92.92 60.46 75.36 71.06 80.22 83.88 91.23 90.76 91.71
ICIF [17] 79.86 88.80 86.85 90.84 61.38 76.07 73.31 91.94 78.27 87.81 84.06 91.96
TFI-GR [48] 83.03 90.73 89.53 91.95 71.38 83.30 81.16 85.56 88.55 93.92 92.65 95.24
DMINet [20] 83.14 90.79 88.89 92.78 68.20 81.10 76.82 85.88 79.53 88.60 91.24 86.10
USSFC-Net [49] 82.54 90.44 91.82 89.09 65.75 79.34 81.52 77.27 83.92 91.26 93.20 89.40
M-Swin * [51] 83.57 91.05 90.05 92.07 -   -   -   - 86.09 92.94 91.37 94.56
ELGC-Net [50] 83.52 91.02 89.78 92.30 65.34 79.03 77.26 80.89 75.56 86.08 82.15 90.40
MDENet * [33] 83.63 91.0992.68 89.55 -   -   -   - 87.74 93.47 95.49 91.54
@@
V-B Implementation Details
Our experiments are conducted on a single NVIDIA GeForce RTX 3090 GPU. The AdamW [52] is employed, with a learning rate of 0.0001 and a weight decay parameter of 0.01. Cross entropy loss is used as our loss function, and the maximum number of iterations is set to 40,000. For data augmentation, RandomFlip and PhotoMetricDistortion are applied. The constant $g$ in the ASFF block is set to 0.1, the decay rate $\gamma$ is set to 0.0001, and $\lambda_{0}$ is set to 1. To evaluate the performance of our experiments, we adopt four widely used metrics: recall (Rec), precision (Pre), F1-score (F1) and intersection over union (IoU), where IoU and F1 are the primary metrics, while the others are auxiliary metrics. The following are formulas for defining the evaluation metrics as mentioned above.